OpenStudy (anonymous):

If x≠2, and x≠-2, then what does (x^3+3x^2-10x)/(2x^2-8) equal?

hartnn (hartnn):

can u factorise the cubic expression in the numerator?

OpenStudy (anonymous):

x(x+5)(x-2)

hartnn (hartnn):

yes and denominator?

OpenStudy (anonymous):

2(x^2-4)

hartnn (hartnn):

which is? use a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

if x≠2, and x≠-2

OpenStudy (anonymous):

i dont get this part

hartnn (hartnn):

since x not =2 --->(x-2) not=0 only then u can cancel out (x-2) from numerator and denominator

OpenStudy (anonymous):

x(x+5)/2(x^2-4)

OpenStudy (anonymous):

and x≠-2

hartnn (hartnn):

to cancel out x-2 from denominator,first factor the denominator as 2(x^2-4) as 2(x+2)(x-2) what is true for x-2,also applies to x+2

OpenStudy (anonymous):

so it's x(x+5)/2(x+2)

hartnn (hartnn):

yup it is :)

OpenStudy (anonymous):

thanx man!